OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (3, -3) with 0° ≤ θ < 360°. choices: (3 , 315°), (-3 , 135°) (3 , 225°), (-3 , 45°) (3 , 45°), (-3 , 225°) (3 , 135°), (-3 , 315°)

jimthompson5910 (jim_thompson5910):

Is the given point (3, -3) in polar form? or cartesian form?

OpenStudy (anonymous):

polar

jimthompson5910 (jim_thompson5910):

ok so I'm guessing that (3, -3) means r = 3 and theta = -3 radians right?

jimthompson5910 (jim_thompson5910):

or I guess theta = -3 degrees not exactly sure

OpenStudy (anonymous):

it would be degrees

jimthompson5910 (jim_thompson5910):

ok well -3 degrees is coterminal to 360 - 3 = 357 degrees

OpenStudy (anonymous):

wait but thats not on there.. lol

jimthompson5910 (jim_thompson5910):

hmm let me think for a sec

OpenStudy (anonymous):

okay haha

jimthompson5910 (jim_thompson5910):

something seems to be missing a symbol maybe can you post a screenshot of the entire problem please?

OpenStudy (anonymous):

i have no idea how to screen shot lol

jimthompson5910 (jim_thompson5910):

above the insert key is the print screen key

jimthompson5910 (jim_thompson5910):

on your keyboard

OpenStudy (anonymous):

but i copy and pasted it word for word and i checked and no symbols are missing

jimthompson5910 (jim_thompson5910):

hit that button then open up MS paint paste it in save it, post it

OpenStudy (anonymous):

oh my gosh i feel like an idiot.. there is something missing

OpenStudy (anonymous):

im sorry let me fix it

jimthompson5910 (jim_thompson5910):

that's ok, and the site is going verrry slow for me lol, so fair warning

jimthompson5910 (jim_thompson5910):

hopefully you figured out how to take a screenshot

OpenStudy (anonymous):

\[(3 \sqrt{2,} 315°), (-3\sqrt{2} , 135°)\] \[ (3\sqrt{2} , 225°), (-3\sqrt{2} , 45°) \] \[ (3\sqrt{2} , 45°), (-3 \sqrt{2}, 225°)\] \[ (3\sqrt{2} , 135°), (-3 , 315°)\]

jimthompson5910 (jim_thompson5910):

oh ok, those are the answer choices?

OpenStudy (anonymous):

oh the last one is \[ (3 \sqrt{2}, 135°), (-3\sqrt{2} , 315°)\]

OpenStudy (anonymous):

yes, my bad

jimthompson5910 (jim_thompson5910):

alright, that's ok and thx

OpenStudy (anonymous):

does that make more sense?

jimthompson5910 (jim_thompson5910):

if the given point is in cartesian form, then r = sqrt(x^2 + y^2) r = sqrt(3^2 + (-3)^2) r = sqrt(9 + 9) r = sqrt(18) r = 3*sqrt(2)

jimthompson5910 (jim_thompson5910):

the angle is theta = arctan(y/x) theta = arctan(-3/3) theta = arctan(-1) theta = -45 which is coterminal to 360 + (-45) = 315 degrees

jimthompson5910 (jim_thompson5910):

now if you add 180 degrees to this, you basically make r negative (because you're going in reverse now) so 315 + 180 = 495 which is coterminal to 495 - 360 = 135 degrees

OpenStudy (anonymous):

oh so it's the first!

jimthompson5910 (jim_thompson5910):

so yep, it's A

OpenStudy (anonymous):

thank you so much i really appreciate the help!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

oh wait i lied.. lol im sorry! hold on let me fix the choices!